OpenStudy (anonymous):

Harris has two pipes to fill his aquarium. Pipe A can fill the aquarium at the rate of 4 gallons per min. Pipe B can fill the aquarium at the rate 840 liters per hour. {1 gallon= 3.70 liters} Part A. Which pipe will fill the aquarium faster? Justify your answer by comparing the rates after converting to the same units. Part B. If the total volume of the aquarium was 500 gallons and you checked on how full it was every 0.25 hours,describe appropiate scales on the x-axis and y axis to graph the rate the faster pipe fills the aquarium

hero (hero):

Convert liters per hour to gallons per minute: \[840 \frac{\text{L}}{\text{hr}} \times \frac{\text{gal}}{3.70 \space\text{L}} \times \frac{\text{hr}}{60 \space\text{min}}\]

OpenStudy (anonymous):

Okay!

hero (hero):

Remember to cancel the appropriate units: \[840 \frac{\cancel{\text{L}}}{\cancel{\text{hr}}} \times \frac{\text{gal}}{3.70 \space\cancel{\text{L}}} \times \frac{\cancel{\text{hr}}}{60 \space\text{min}}\] Which leaves: \[\frac{840}{(3.70)(60)} \frac{\text{gal}}{\text{min}}\]

hero (hero):

Simplify the result

OpenStudy (anonymous):

Wow! okay thanks so much!! You really helped me out.

hero (hero):

You're Welcome

OpenStudy (anonymous):

:0)